OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

OpenStudy (uri):

m=(y2-y1)/(x2-x1) m=-6-8/-3-(-4) m=-14/1

OpenStudy (anonymous):

Ok so m=(y2-y1)/(x2-x1) is the equation used to find the slope right?

OpenStudy (anonymous):

y = -0.25x + -4

OpenStudy (uri):

yes @massify the formula :3